OpenStudy (anonymous):

Water freezes at thirty-two degrees Fahrenheit, or zero degrees Celsius. Use temperature conversion formulas to show that these two measures are indeed the same temperature.

OpenStudy (anonymous):

\[F = C * \frac{9}{5} + 32\] since our C = 0 \[0 * \frac{9}{5} + 32 = 0 + 32 =32=F\]

OpenStudy (anonymous):

32 F = 0 C

OpenStudy (anonymous):

thankyou :) @hobbs978

OpenStudy (anonymous):

no problem :)